Joe plans to make ropes and matching hats from a bolt of material. The fabric he plans to use is 250 yards long. She will need 2 yards of material for each rope and 1/6 of a yard for each hat She also plans to put decorations on the rope and will need 1 yard of cord for each rope . The hats will not have any fingers . She has 100 yards of cord to use for the decorations.
1. Write two equations to use to solve for how many hats and ropes she can make from the supplies on hand. Let r equal the number of ropes and made and h equal the number of hats made

2. Try to solve this system of equations If you get a solution state what it is. If not describe what the result mean.
: Joe plans to make ropes and matching hats from a bolt of material. The fabric he plans to use is 250 yards long. She will need 2 yards of material for each rope and 1/6 of a yard for each hat She also plans to put decorations on the rope and will need 1 yard of cord for each rope . The hats will not have any fingers . She has 100 yards of cord to use for the decorations.
1. Write two equations to use to solve for how many hats and ropes she can make from the supplies on hand. Let r equal the number of ropes and made and h equal the number of hats made

2. Try to solve this system of equations If you get a solution state what it is. If not describe what the result mean.

Answer by stanbon(18984)

Joe plans to make ropes and matching hats from a bolt of material.
The fabric he plans to use is 250 yards long.
He will need 2 yards of material for each rope and 1/6 of a yard for each hat He also plans to put decorations on the rope and will need 1 yard of cord for each rope .
The hats will not have any fingers .
He has 100 yards of cord to use for the decorations.
-----------------------
1. Write two equations to use to solve for how many hats and ropes he can make from the supplies on hand. Let r equal the number of ropes and made and h equal the number of hats made.
Fabric inequality: 2r + (1/6)h <= 250
Cord inequality: r <= 100
-------------------------

2. Try to solve this system of equations If you get a solution state what it is. If not describe what the result mean.
Solve the EQUALITY;
2*100 + (1/6)h = 250
(1/6)h = 50
h = 300
So, when # of ropes is 100, # of hats is 300.
-----------
Cheers,
Stan H.